pooja195 (pooja195):

Solve using substitution. I know some of this but I always end up screwing up the final answer it would be awesome if someone can guide me through it :) 2x – 3y = –2 4x + y = 24

hartnn (hartnn):

take the 2nd equation try to isolate y

OpenStudy (anonymous):

Rearrange the second equation to get an expression for y

pooja195 (pooja195):

i would subtract y right?

OpenStudy (anonymous):

Subtract 4x

hartnn (hartnn):

4x+y = 24 subtract 4x from both sides

hartnn (hartnn):

4x-4x +y = 24-4x

OpenStudy (anonymous):

will it be 4x = 24 - y

pooja195 (pooja195):

right so i end up with y= -4x + 24

OpenStudy (anonymous):

Substitute this expression for y into the first equation

OpenStudy (anonymous):

ohh yes poo carry on

pooja195 (pooja195):

so after that i would sub for the first equation and get 2(-4x+24)-3y=-2?

OpenStudy (anonymous):

No. Substitute for y, giving 2x -3(-4x+24)=-2. The solve for x.

OpenStudy (anonymous):

Remember, by rearranging the second equation, you got an expression for y, not x,

pooja195 (pooja195):

Whoops my bad sorry um.... so i would end up with 2x+12x-72=-2?

OpenStudy (anonymous):

Terrific. Now solve for x.

pooja195 (pooja195):

14x-72=-2? add 72 14x=70?

OpenStudy (anonymous):

Yup. Keep going.

pooja195 (pooja195):

x=5

OpenStudy (anonymous):

Right on. Now you know the solution for x. Choose either of the two equations from the problem, substitute 5 for x, and solve for y.

pooja195 (pooja195):

4(5) + y = 24 20+y=24?

pooja195 (pooja195):

y=4

OpenStudy (anonymous):

Well done. Congrats.

pooja195 (pooja195):

Yay!!! thanks<3

OpenStudy (anonymous):

You're welcome.